Copy of If Z is a standard normal random variable, find P(-1.7 <Z < 2.6). 0.4217 0.9507 -0.9508 0.9953 0.0446
Answer #1

OPTION "B" 0.9507 IS CORRECT

TO OBTAIN THE ANSWER WE NEED TO FOLLOW SEVERAL STEPS THAT ARE GIVEN BELOW

Step 1: Sketch the curve. The probability that -1.7<7<2.6 is equal to the blue area under the curve. 0.50 0.45 0.40 0.35 0.30

0.09 Z 0.0 0.1 02 03 04 05 06 0.7 08 09 1.0 11 12 13 14 1.5 16 17 1.8 19 20 21 0.00 0.01 0.02 0.03 0.04 10.05 0.06 0.07 0.08

We see that P(Z < 2.6 ) = 0.9953. Step 4: P(Z < -1.7 ) can be found by using the following fomula. P(Z <-a)=1 - P(Z <a) After

Z Z 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 2.0 2.1 2.2 23 2.4 2.5 2.6 0.00 0.01 0.02

We see that P ( 2<1.7) = 0.9554 so, P(Z < -1.7) = 1 - P(Z <1.7)=1 -0.9554 = 0.0446 At the end we have: P(-1.7<7 < 2.6 ) = 0.9

IN THIS WAY WE FIND THE VALUE OF P(-1.7<Z<2.6) = 0.9507

_______________________________________________________________________________________